# BlockLLM: Multi-tenant Finer-grained Serving for Large Language Models

### v1

@@ 7 Evaluation
We evaluate BlockLLM using testbed experiments. We seek to answer two main questions:
@@ 7.1 Setup
Cluster. Our cluster has four servers. Two are equipped with two A100 GPUs with 80GB memory, and the other two have four A100 GPUs with 80GB memory. The servers are interconnected with 100Gbps network.
@@
Workload. We generate a synthetic workload trace (Figure 15) to evaluate BlockLLM’s performance. We use uniform distribution to generate a mean rate for each application. The mean rate is used as the weight to calculate the number of requests for each application, so some applications are more popular. Then, the arrival time of each request in one application is generated using a Poisson process with the given mean rate. The trace is 20 minutes, and there are 400 requests. The generation approach is similar to [34].
@@
Baselines. We benchmark BlockLLM with two other solutions.
@@
Metrics. We evaluate BlockLLM with metrics, including latency of serving one request, throughput (number of tokens per second), communication costs (percentage of time spent on data transferring), GPU utilization (SM efficiency), and device memory consumption.
BlockLLM’s configuration. In this evaluation, we apply speculation to the top 10% bottlenecked block instances sorted by the time of completing their request queues. BlockLLM’s scheduler checks redundant KV cache every one minute. We set 0.95 as the cosine similarity threshold for determining whether BlockLLM’s surrogate prediction is accurate and 0.98 as the threshold for determining equivalence. The maximum sequence length allowed is 1024. BlockLLM trains the stitching block with the MMLU [14].
@@ 7.2 Overall Performance
We first compare the overall performance of BlockLLM. The default number of applications is 20.
@@
Latency and throughput. Figure 15 depicts the CDF of the latency of completing a request in BlockLLM. BlockLLM’s median latency is 3.34, comparable with PM (3.39) and PS (3.87) provisioning. The reduction of 95%ile latency is more significant, 33.5% and 23.4% compared with PM and PS. The throughput of BlockLLM is 1.71x of PM and 1.46x of PS provisioning (Figure 17). Since BlockLLM allows larger batch sizes with higher efficiency, the tail latency is improved significantly. Though PS can serve various applications within one model, the cost of one inference process grows due to branching.
@@
Figure 19: Performance changes when the the number of applications grows.
GPU utilization. We also measure overall GPU utilization indicated by SM efficiency and memory consumption in the cluster. We monitor the end-to-end serving process, and the average GPU utilization is improved by 20.1% and 4.8% compared with PM and PS provisioning. Figure 17 depicts the GPU utilization change over the serving process. With the help of adaptive serving, BlockLLM efficiently dispatches requests under the existing deployment status to avoid model frequent model loading and unloading. As for memory consumption, we measure the memory consumption of model parameters and request-related matrices (including input, intermediate activations, output, and KV cache). In the best case, BlockLLM takes 16.1% less space on model parameters and 24.1% more space on request-related matrices, indicating more data are being processed.
@@ 7.3 Effectiveness of BlockLLM’s Design
We then evaluate the effectiveness of BlockLLM’s key designs using an ablation study.
Figure 20: Cosine similarity of requests using adaptive serving and predetermined block chains.
@@
Locality-aware block placement. We compare the communication costs between BlockLLM’s locality-aware placement and the widely adopted fragmentation-minimized placement. Figure 23 shows the average performance change of using fragmentation-minimized placement. The median and 95%ile latency is increased by 12.6% and 18.2%. The communication costs of one request sum up all the transfer costs, therefore presenting a significant inflation of 73.4%. The locality-aware placement has reduced 72.3% inter-server communications compared with the fragmentation-minimized placement strategy.
@@ 9 Related Work
System for LLM models. Model-internal optimization including FasterTransformer [2], PagedAttention [18], FlashAttention [8] and FlexGen [35] are complimentary to BlockLLM’s contributions. Orca [42] introduces iteration-level scheduling to handle the sequence length difference among multiple requests in the batch. SpotServe [27] realizes LLM serving on preemptible instances with communication cost-minimized migration plan and stateful recovery. ORCA and SpotServe can be integrated with BlockLLM’s scheduler and agent. PetS [45] expresses four representative PE techniques in a unified representation so that the requests from four applications can be executed simultaneously, similar to Parameter Sharing evaluated in 7.
Auto-regressive. The inherent auto-regressive nature of LLMs limits their serving efficiency, making parallel decoding a primary focus of recent research. SpecInfer [26] and [19] circumvents this constraint through speculative execution. It employs smaller LLM models to predict multiple tokens in advance. [29, 11] further improve the solution by removing the need for surrogate models. Instead, they use the existing model together with a few look-ahead tokens to make predictions. All these strategies aim to generate multiple preliminary drafts, thereby facilitating earlier production of the subsequent token. BlockLLM, however, implements speculation with a more refined approach. BlockLLM applies speculative execution at the block-level in a selective manner, optimizing the generation process at the individual token level.

### v2

@@ 7 Evaluation
We evaluate BlockLLM using testbed experiments. We seek to answer three main questions:
  • What impact does BlockLLM has on accuracy? (7.2)
@@ 7.1 Setup
Cluster. Our cluster has four servers, each of which is equipped with A100 GPUs with 80GB memory. Two servers have two GPUs, and the other two have four GPUs. The servers are interconnected with 100 Gbps links.
@@
Workload. We generate both production and synthetic workload traces (Figure 18) to evaluate BlockLLM’s performance. For synthetic traces, we use uniform distribution to determine the mean rate for each application. The mean rate serves as the weight to calculate the number of requests per application, creating varying popularity among applications. The arrival time of each request in one application is generated using a Poisson process with the given mean rate. The trace spans 20 minutes and includes 400 requests, following a generation approach similar to [43]. For production traces, we use timing information from a Twitter trace collected over a month-long period [3]. We set the minimum QPS to 1 and maximum QPS to 45. Previous work has shown this trace accurately reflects realistic inference workloads, exibiting diurnal patterns and unexpected spikes [54].
Figure 16: Accuracy and request arrival rate changes.
@@
Baselines. We benchmark BlockLLM with two baselines.
@@
Metrics. We evaluate BlockLLM with several metrics, including latency of serving one request, throughput (number of tokens per second), communication costs (percentage of time spent on data transferring), GPU utilization (SM efficiency), device memory consumption, and end-to-end accuracy.
BlockLLM’s configuration. In this evaluation, we apply speculative execution to the top 10% bottlenecked block instances sorted by the time of completing their request queues. BlockLLM’s scheduler checks redundant KV cache every one minute. We set 0.95 as the cosine similarity threshold for determining whether BlockLLM’s surrogate prediction is accurate and 0.98 as the threshold for determining equivalence. The maximum sequence length allowed is 1024. BlockLLM trains the stitching block with the MMLU [19].
@@ 7.2 Overall Performance
We first evaluate the overall performance of BlockLLM. The default number of applications is 20.
@@
Accuracy. We use the production Twitter traces to study how accuracy reacts to changing request arrival rate. Figure 16 illustrates the variation in accuracy over time under different arrival rates. The models evaluated include LLaMA 1.1B, 7B, and 13B. Incoming requests prioritize the highest-accuracy blocks (13B) unless resources are unavailable. The average accuracy achieved is 42.1%, which is 3.5% higher than the closest option (7B). Higher request arrival rates cause BlockLLM to redirect traffic to faster but less accurate blocks (1.1B) more frequently, resulting in a gradual decline in accuracy.
Latency and throughput. Figure 18 depicts the CDF of the latency of completing a request in BlockLLM. BlockLLM’s median latency is 3.34 min, comparable with PM (3.39 min) and PS (3.87 min). The reduction of 95%ile latency is more significant, 33.5% and 23.4% compared with PM and PS. The throughput of BlockLLM is 1.71x of PM and 1.46x of PS (Figure 20). By decomposing models into more granular blocks, BlockLLM enhances the efficiency of processing larger batch sizes. This approach significantly reduces tail latency, especially under high request rates. Though PS can serve various applications within one model, the cost of one inference process grows due to branching.
@@
Figure 22: Performance changes when the number of applications grows.
GPU utilization. We also measure overall GPU utilization indicated by SM efficiency and memory consumption in the cluster. We monitor the end-to-end serving process, and the average GPU utilization is improved by 20.1% and 4.8% compared with PM and PS. Figure 20 depicts the GPU utilization change over the serving process. With the help of adaptive serving, BlockLLM efficiently dispatches requests under the existing deployment status to avoid frequent model loading and unloading. We measure the memory consumption of model parameters and request-related matrices (including input, intermediate activations, output, and KV cache). In the best case, BlockLLM takes 16.1% less space on model parameters and 24.1% more space on request-related matrices, indicating more data are being processed.
@@ 7.3 Effectiveness of BlockLLM’s Design
We then evaluate the effectiveness of BlockLLM’s key designs via an ablation study.
Figure 23: Cosine similarity of output vocabulary using adaptive serving and predetermined block chains.
@@
Locality-aware block placement. We compare the communication costs between BlockLLM’s locality-aware placement and the widely adopted fragmentation-minimized (frag-min) placement. Figure 26 shows the average performance change of using the frag-min placement. The median and 95%ile latency is increased by 12.6% and 18.2%. The communication costs of one request sum up all the transfer costs, therefore presenting a significant inflation of 73.4%. The locality-aware placement has reduced 72.3% inter-server communications compared with the frag-min placement strategy.
@@ 9 Related Work
System for LLM models. Model-internal optimization including FasterTransformer [2], PagedAttention [25], FlashAttention [13] and FlexGen [44] are complimentary to BlockLLM’s contributions. Orca [51] introduces iteration-level scheduling to handle the sequence length difference among multiple requests. SpotServe [35] enables LLM serving on preemptible instances with minimized communication costs and stateful recovery. PetS [55] unifies four representative PE techniques, allowing simultaneous execution of requests from different applications, similar to Parameter Sharing in 7.  [22, 5, 42] investigate accuracy-scaling to adapt to fluctuating workloads.
Auto-regressive. The inherent auto-regressive nature of LLMs limits their efficiency, making parallel decoding a key research focus. SpecInfer [34] and [26] address the issue with speculative execution, using smaller models to predict multiple tokens in advance. [37, 16] improve this by eliminating surrogate models, using the existing model with a few look-ahead tokens. These strategies generate multiple drafts to speed up token production. On the other hand, BlockLLM applies speculative execution at the block-level in a selective manner, optimizing the generation process at the individual token level.